OpenStudy (anonymous):

cos(x+pi/2) = -sin x help! i need somebody.. help! not just anybody

hartnn (hartnn):

Apply cos (A+B) formula know what it is ?

hartnn (hartnn):

\(\cos(A+B) = \cos A\cos B - \sin A \sin B \)

hartnn (hartnn):

A = x B =pi/2

OpenStudy (anonymous):

so, cos x cos pi/2 - sin x sin pi/2

hartnn (hartnn):

yes what are cos pi/2 = .. sin pi/2 = .. ?

OpenStudy (anonymous):

do they equal each other?

hartnn (hartnn):

no... cos pi/2 = 0 sin pi/2 = 1 do you have unit circle with you ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

then look for pi/2 angle in it. what are the co-ordinates when angle is pi/2 radians or 90 degrees

OpenStudy (anonymous):

(0,1)

hartnn (hartnn):

correct, those are read as \((\cos \theta , \sin \theta)\) and since \(\theta = \pi/2 \) we get \((\cos \pi/2 , \sin \pi/2 )=(0,1)\)

hartnn (hartnn):

thats why cos pi/2 = 0 and sin pi/2 = 1 now plug these in your equation

OpenStudy (anonymous):

cos x 0 - sin x 1 ?

hartnn (hartnn):

correct! simplify it

OpenStudy (anonymous):

- sin x

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

ahh! you're a life saver!

hartnn (hartnn):

^_^

OpenStudy (anonymous):

thank you sososososososososososo much!

hartnn (hartnn):

welcome sososososososososososo much!

OpenStudy (anonymous):

:)